can you please send me some practice work on algebra? i'm having problem with problem like 2x + 4y= 8
6x + 12y= 24
Then it would ask question like, what is the value of Y? I do run into problem like 3x + 4y= 12, and 6x + 14y= 30, so what is the y + 10=
can you email how do problem like this

Found 2 solutions by askme, BLAZETHEKILLER:
Answer by askme(9)

your first question has the same line, so we will solve your second.
since you have two variables (two unknowns - x and y), they are giving you two equations so that you can solve for each variable by substitution.


step 1: write equations in linear format y = mx + b
so for 3x + 4y = 12, we write it as 4y = -3x + 12.
for 6x + 14y = 30, we write it as 14y = -6x + 30


step 2: simplify your equations.
for the equation 4y = -3x + 12, divide the whole equation by 4 so that the coefficient of y will equal 1.
4y/4 = -3x/4 + 12/4
after dividing, this will be written as y = -3x/4 + 3
2nd equation: 14y = -6x + 30 will be divided by 14 and will be rewritten as y = -3x/7 + 15/7


step 3: substitute y in the first equation since there are two variables.
since y = -3x/7 + 15/7, when we combine your 2nd equation with your 1st, we write it as -3x/7 + 15/7 = -3x/4 + 3.


step 4: bring x's to one side of the equation and your constants to one side.
-3x/7 + 3x/4 = 3 - 15/7


step 5: find the lowest common denominator to get rid of the fractions. thus, we will multiply the whole equation of -3x/7 + 3x/4 = 3 - 15/7 by 28 and get:
-12x + 21x = 3(28) - 60
9x = 84 - 60
9x = 24
x = 24/9 = 8/3



step 6: subsitute x into your original equation. we will use the 1st one 3x + 4y = 12
3(8/3) + 4y = 12
(the 3's cancel)
8 + 4y = 12
4y = 12 - 8
4y = 4
y = 1



Thus, x = 8/3 and y = 1

Answer by BLAZETHEKILLER(3)

ITS HARD HEHEHEHEHHEHEHE